if one-half of the supplement of an angel is six degrees less than twice the complement the measure of the angel is ______ degrees.
I did not really know how to write this equation. Please help me.
1/2 (180-x)=6-2(90-x)

Answer by rfer(10417)

.5(180-x)=2(90-x)-6
90-.5x=180-2x-6
1.5x=90-6
1.5x=84
x=56
------------
180-56=124/2=62
90-56=34*2=68-6=62
--------------
Your basic setup was good except that you need to subtract the 6 from 2(90-x)
reverse what you had.
Good luck
Bob